OpenStudy (anonymous):

Algebra 2 . Substitution. mutliple choice!

OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 2x – y = –4 5x – 6y = 12 5(–2x – 4) – 6y = 12 5(2x + 4) – 6y = 12 5x – 6(–2x – 4) = 12 5x – 6(2x + 4) = 12

OpenStudy (anonymous):

can you solve \(2x-y=-4\) for \(y\)?

OpenStudy (anonymous):

would you subtract 2x from both sides

OpenStudy (anonymous):

yes you could do that you would get \[-y=-2x-4\] but you have one more step to get \(y\)

OpenStudy (anonymous):

i meant like. 2x-y=-4 -2x -2 y=-2?

OpenStudy (anonymous):

no like this 2x - y = -4 -2x -2x -y = -2x - 4

OpenStudy (anonymous):

if you subtract \(2x\) from both sides you get \(-2x-4\) on the right

OpenStudy (anonymous):

oh okay. so whats next

OpenStudy (anonymous):

hope that is clear, you do not subtract 2x from the left and 2 from the right next step is to solve for \(y\) you have \(-y=-2x-4\) and you want \(y\) not \(-y\) so multiply both sides by \(-1\) which is the same as changing all the signs

OpenStudy (anonymous):

what do you get ?

OpenStudy (anonymous):

y=2x+4?

OpenStudy (anonymous):

exactly!

OpenStudy (anonymous):

now for the last step replace \(y\) in the equation \(5x – 6y = 12 \) by \(2x+4\) use parentheses

OpenStudy (anonymous):

is that step clear? it should look exactly like on of your possible answers above

OpenStudy (anonymous):

5x-6(2x+4)=12?

OpenStudy (anonymous):

you got it

OpenStudy (anonymous):

thankyou very much :) so it's the last one?!

OpenStudy (anonymous):

yes it is the last one